The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 45 June 25, 2000

Contents

Abstracts of 11 refereed papers 1
Meeting announcements  9

Abstracts of Refereed Papers

Mid-Infrared imaging and spectrophotometry of N 66 in the SMC with ISOCAM

A. Contursi^{1,2,3}, J. Lequeux^{2}, D. Cesarsky^{4,5}, F. Boulanger^{4}, M. Rubio^{6}, M. Hanus^{2}, M.

Sauvage^{1}, D. Tran^{5}, A. Bosma^{7}, S. Madden^{1}, L. Vigroux^{1}

1 SAp/DAPNIA/DSM, CEA-Saclay, F-91191 Gif sur Yvette CEDEX, France 2 DEMIRM, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 3 IPAC, Caltech, MS 100–22, Pasadena, CA 91125, USA 4 Institut d’Astrophysique Spatiale, Bat. 121, Université Paris XI, 91450 Orsay CEDEX, France 5 Max–Planck Institut für extraterrestrische Physik, Postfach 1603, D-85740, Garching, Germany 6 Departamento de Astronomia, Universidad de Chile, Casilla 360D, Santiago, Chile 7 Observatoire de Marseille, 2 Place le Verrier, F-13248 Marseille CEDEX, France

We present observations with the mid–infrared camera ISOCAM on board the Infrared Space Observatory of the major star-forming region N 66 in the Small Magellanic Cloud (SMC) and of its surroundings. These observations were performed with broad filters and Circular Variable Filters giving a spectral resolution of about 40. In addition, CO(2-1) data are presented, allowing us to identify and study how hot dust relates with the different phases of the Interstellar Medium (ISM) present in N66. The spectra are dominated by the strong emission of fine-structure line. Monochromatic maps have been made in the [Ne iii] 15.6 μm and [S iv] 10.5 μm line. There are significant differences between their distributions, due to the effects of density and of shocks. Aromatic Infrared Bands (AIBs) are seen at various places in the field but they are generally faint. They exhibit a variety of shapes and relative intensities, suggesting that a diversity of carbonaceous materials are present. Silicate emission is also clearly visible in the central condensation and in a few others and emission from hot small grains (Very Small Grains, VSGs) longward of 10 μm is present in the whole region. All these dust components are heated by the very strong far-UV radiation of the many young, massive stars contained in the region. The interstellar radiation field (ISRF) at 1600 Å is ≥ 10^{5} times the ISRF of the solar neighborhood in the peaks of mid-infrared emission. The relative contributions of these components (AIB carriers, VSGs and silicate grains) to the mid-infrared spectra seem to depend on the intensity and the hardness of the far-UV field. In general the 15/6.75 μm intensity ratio is higher than in relatively quiescent galactic regions (Cesarsky et al. 1996b, Abergel et al. 1996) but it is not as high as expected for a linear increase with ISRF. We interpret this behavior as due to the destruction of both AIBs carriers and VSGs in a very high ISRF. Finally several stars have been detected at 6.75 μm. Two are red supergiants; the other stars are blue and the IR emission is due to circumstellar matter or to interstellar matter heated by the star.

Accepted by: Astronomy & Astrophysics

For preprints, contact contursi@ipac.caltech.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0006185

Radio Properties of the Supernova Remnant N157B

J.S. Lazendic^{1}, J.R. Dickel^{2}, R.F. Haynes^{3}, P.A. Jones^{1}, and G.L. White^{1}

1 Centre for Astronomy, University of Western Sydney, PO Box 10, Kingswood NSW 2747, Australia 2 Astronomy Department, University of Illinois at Urbana-Champaign, 1002 West Green Street, Urbana IL 61801, USA 3 Australia Telescope National Facility, PO Box 76, Epping NSW 1710, Australia

A new investigation of the supernova remnant (SNR) N157B was carried out with the Australia Telescope Compact Array. Radio continuum images of the entire 30 Doradus region have been made at 3.5 and 6 cm wavelength with a resolution of 2”. These data allow a high resolution study of the spectral index distribution and polarization properties of both N157B and the nearby 30 Doradus nebula (the latter will be reported in a subsequent paper). N157B is an extended Crab-type SNR which may be beginning the transition to a composite remnant. There is little apparent fine structure and the brightest radio region is several parsecs from the probable position of the X-ray pulsar. The SNR has a radio spectral index of -0.19 and is significantly polarized at 3.5 cm but not at longer wavelengths.

Accepted by: The Astrophysical Journal

For preprints, contact Jasmina.Lazendic@atnf.csiro.au

Dredge-up Effects in Galactic and Magellanic Planetary Nebulae

R.D.D. Costa^{1}, J.A. de Freitas Pacheco^{2}, T.P. Idiart^{1}

1Instituto Astronômico e Geofísico, Av. Miguel Stefano 4200, 04301-904, S.Paulo, Brazil 2Observatoire de la Côte d’Azur, B.P. 4229, 06304 Nice, France

Chemical abundances are reported for 23 planetary nebulae in the Small Magellanic Cloud. Average abundances agree with values found for H ii regions, suggesting that progenitors of most of these objects have been formed in the last 1–2 Gyr. The N/O vs. O/H anticorrelation is confirmed for Magellanic and galactic planetaries, and a more robust result is obtained when objects of similar ages (or masses) are grouped together. For a given class, such a negative correlation suggests that the surface enrichment due to dredge-up episodes is more efficient at lower metallicities, in agreement with recent model computations.

Accepted by: Astronomy & Astrophysics Supplement For preprints, contact R.D.D. Costa roberto@orion.iagusp.usp.br

The Detection of Wind Variability in Magellanic Cloud O Stars

D. Massa^{1}, A.W. Fullerton^{2,3}, J.B. Hutchings^{4}, D.C. Morton^{4}, G. Sonneborn^{5}, A.J. Willis^{6}, L. Bianchi^{3}, K.R. Brownsberger^{7}, P.A. Crowther^{7}, T.P. Snow^{7} and D.G. York^{8}

1Raytheon ITSS, Greenbelt, USA 2U. Victoria, Victoria, Canada 3JHU, Baltimore, USA 4DAO, Herzberg Inst. Ap., Victoria, Canada 5NASA’s GSFC, Greenbelt, USA 6UCL, London, UK 7U. Colorado, Boulder, USA 8U. Chicago, Chicago, USA

We present Far Ultraviolet Explorer (FUSE) spectra for three Magellanic Cloud O stars (Sk 80, Sk −67^{∘}05 and Sk −67^{∘}111) with repeated observations. The data demonstrate the capabilities of FUSE to perform time-resolved spectroscopy on extragalactic stars. The wavelength coverage of FUSE provides access to resonance lines due to less abundant species, such as sulfur, which are unsaturated in O supergiants. This allows us to examine wind variability at all velocities in resonance lines for stars with higher mass loss rates than can be studied at longer (λ ≥ 1150Å) wavelengths. The FUSE wavelength range also includes resonance lines from ions which bracket the expected dominant ionization stage of the wind. Our observations span 1–4 months with several densely sampled intervals of 10 hours or more. These observations reveal wind variability in all of the program stars and distinctive differences in the ionization structure and time scales of the variability. Sk −67^{∘}111 demonstrates significant wind variability on a time scale less than 10 hours and the coolest O star (Sk −67^{∘}05) exhibits the largest variations in O vi.

Accepted by: ApJ Letters FUSE special edition

For preprints, contact derck.massa@gsfc.nasa.gov
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0005434

Theoretical Models for Classical Cepheids: VII. Metallicity effects on the Cepheid distance scale

F. Caputo^{1}, M. Marconi^{2} and I. Musella^{2}

1 Astron. Obs. of Roma, Italy 2 Astron. Obs. of Capodimonte, Italy

We use theoretical Period-Luminosity and Period-Luminosity-Color relations in the VI passbands, as based on nonlinear, nonlocal and time-dependent convective pulsating models, to predict the reddening and true distance modulus of distant Cepheids observed with the Hubble Space Telescope. By relying on the pulsating models with metal content Z=0.008, we find that the theoretical predictions agree to the values obtained by the Extragalactic Distance Scale Key Project on the basis of empirical Period-Luminosity relations referenced to LMC variables. In the meantime, from the theoretical relations with Z=0.004 and 0.02 we find that the predicted E(B–V) and μ_{0} decrease as the adopted metal content increases. This suggests a metallicity correction to LMC-based distances as given by about −0.27 mag/dex. Such a theoretical correction appears supported by an existing, although weak, correlation between the Cepheid distance and the [O/H] metallicity of galaxies within a given group or cluster, as well as by a similar correlation between the H_{0} estimate and the [O/H] metallicity of the galaxies which calibrate the SNIa luminosity. On the contrary, the metallicity correction earlier suggested on empirical grounds seem to be excluded. Eventually we suggest that the average value ⟨H_{0}⟩ ∼ 67 km/s/Mpc provided by the Key Project team should increase at least up to ∼ 69 km/s/Mpc.

Accepted by: Astronomy & Astrophysics

For preprints, contact marcella@na.astro.it

Classical Cepheid pulsation models. VI. The Hertzsprung progression

G. Bono^{1}, M. Marconi^{2} and R.F. Stellingwerf^{3}

1 Astron. Obs. of Roma, Italy 2 Astron. Obs. of Capodimonte, Italy 3 SC, 2229 Loma Linda, Los Alamos, USA

We present the results of an extensive theoretical investigation on the pulsation behavior of Bump Cepheids. We constructed several sequences of full amplitude, nonlinear, convective models by adopting a chemical composition typical of Large Magellanic Cloud (LMC) Cepheids (Y=0.25, Z=0.008) and stellar masses ranging from 6.55 to 7.45 solar masses. We find that theoretical light and velocity curves reproduce the HP, and indeed close to the blue edge the bump is located along the descending branch, toward longer periods it crosses at first the luminosity/velocity maximum and then it appears along the rising branch. In particular, we find that the predicted period at the HP center is 11.24 ± 0.46 d and that such a value is in very good agreement with the empirical value estimated by adopting the Fourier parameters of LMC Cepheid light curves. We also find that the models at the HP center are located within the resonance region but not on the 2:1 resonance line. Interestingly enough, the predicted Bump Cepheid masses, based on a Mass-Luminosity (ML) relation which neglects the convective core overshooting, are in good agreement with the empirical masses of Galactic Cepheids estimated by adopting the Baade-Wesselink method. Even by accounting for the metallicity difference between Galactic and LMC Cepheids, this result seems to settle down the long-standing problem of the Bump mass discrepancy. Finally, the dynamical behavior of a cool Bump Cepheid model provides a plain explanation of an ill-understood empirical evidence.

Accepted by: Astronomy & Astrophysics

For preprints, contact bono@coma.mporzio.astro.it

Intermediate-mass star models with different helium and metal contents

G. Bono^{1}, F. Caputo^{1}, S. Cassisi^{2,3}, M. Marconi^{4}, L. Piersanti^{2,5}, and A. Tornambè^{2}

1 Osservatorio Astronomico di Roma, Via Frascati 33, 00040 Monte Porzio Catone, Italy 2 Osservatorio Astronomico di Collurania, Via M. Maggini, 64100 Teramo, Italy 3 Max-Planck-Institut fur Astrophysik, Karl-Schwarzschild-Strasse 1, 85740 Garching bei München, Germany 4 Osservatorio Astronomico di Capodimonte, Via Moiariello 16, 80131 Napoli, Italy 5 Università di Napoli, Dip. di Fisica, 80131 Napoli, Italy

We present a comprehensive theoretical investigation of the evolutionary properties of intermediatemass stars. The evolutionary sequences were computed from the Zero Age Main Sequence up to the central He exhaustion and often up to the phases which precede the carbon ignition or to the reignition of the H-shell which marks the beginning of the thermal pulse phase. The evolutionary tracks were constructed by adopting a wide range of stellar masses (3 ≤ M_{⊙} ≤ 15) and chemical compositions. In order to account for current uncertainties on the He to heavy elements enrichment ratio (Δ Y/Δ Z), the stellar models were computed by adopting at Z=0.02 two different He contents (Y=0.27, 0.289) and at Z=0.04 three different He contents (Y=0.29, 0.34, and 0.37). Moreover, to supply a homogeneous evolutionary scenario which accounts for young Magellanic stellar systems the calculations were also extended toward lower metallicities (Z=0.004, Z=0.01), by adopting different initial He abundances.

We evaluated for both solar (Z=0.02) and super-metal-rich (SMR, Z=0.04) models the transition mass M^{up} between the stellar structures igniting carbon and those which develop a full electron degeneracy inside the carbon-oxygen core. We found that M^{up} is of the order of 7.7 ± 0.5M_{⊙} for solar composition, while for SMR structures an increase in the He content causes a decrease in M^{up}, and indeed it changes from 9.5 ± 0.5M_{⊙} at Y=0.29, to 8.7 ± 0.2M_{⊙} at Y=0.34, and to 7.7 ± 0.2M_{⊙} at Y=0.37. We also show that M^{up} presents a nonlinear behavior with metallicity, and indeed it decreases when moving from Z=0.04 to Z ≈ 0.001 and increases at lower metal contents. This finding confirms the predictions by Cassisi & Castellani (1993) and more recently by Umeda et al. (1999) and suggests that the rate of SNe type Ia depends on the chemical composition of the parent stellar population.

This approach allows us to investigate in detail the evolutionary properties of classical Cepheids. In particular, we find that the range of stellar masses which perform the blue loop during the central He-burning phase narrows when moving toward metal-rich and SMR structures. This evidence and the substantial decrease in the evolutionary time spent by these structures inside the instability strip bring out that the probability to detect long-period Cepheids in SMR stellar systems is substantially smaller than in more metal-poor systems.

Moreover and even more importantly, we find that the time spent by Cepheids along the subsequent crossings of the instability strip also depends on the stellar mass. In fact, our models suggest that low-mass, metal-poor Cepheids spend a substantial portion of their lifetime along the blueward excursion of the blue loop, while at higher masses (M_{⊙} ≥ 8) the time spent along the redward excursion becomes longer. Models at solar chemical composition present an opposite behavior, i.e., the time spent along the redward excursion is longer than the blueward excursion among low-mass Cepheids and vice versa for high-mass Cepheids. Oddly enough, the time spent along the blueward excursion is for models at Z=0.01 longer than the redward excursion over the entire mass range. This suggests a nonlinear dependence of crossing times on metallicity. The time spent along the first crossing of the instability strip is generally negligible with the exception of high-mass, metal-poor stellar structures for which it becomes of the order of 15–20% of the total crossing time.

Accepted by: The Astrophysical Journal For preprints, contact S. Cassisi, cassisi@astrte.te.astro.it

Eclipsing Binaries in the OGLE Variable Star Catalogs. V. Long-Period β Lyrae-type Systems in the Small Magellanic Cloud

and the PLC-β Relation

Slavek M. Rucinski^{1} and Carla Maceroni^{2}

1 David Dunlap Observatory, University of Toronto, P.O.Box 360, Richmond Hill, Ontario, Canada L4C 4Y6 2 Rome Observatory, via Frascati 33, I-00040 Monteporzio C., Italy

Thirty eight long-period (P > 10 days), apparently contact binary stars discovered by the OGLE-II project in the SMC appear to be β Lyrae-type systems with ellipsoidal variations of the cool components dominating over eclipse effects in the systemic light variations and in the total luminosity. A new period–luminosity–color (PLC) relation has been established for these systems; we call it the PLC-β relation, to distinguish it from the Cepheid relation. Two versions of the PLC-β relation – based on the (B − V)_{0} or (V − I)_{0} color indices – have been calibrated for 33 systems with (V − I)_{0} > 0.25 spanning the orbital period range of 11 to 181 days. The relations can provide maximum-light, absolute-magnitude estimates accurate to ϵM_{V} ≃ 0.35 mag within the approximate range −3 < M_{V} < +1. In terms of their number in the SMC, the long-period β Lyrae-type binaries are about 50 times less common than the Cepheids. Nevertheless, their large luminosities coupled with continuous light variations make these binaries very easy to spot in nearby galaxies, so that the PLC-β relation can offer an auxiliary and entirely independent method of distance determination to nearby stellar systems rich in massive stars. The sample of the long-period β Lyrae systems in the SMC analyzed in this paper is currently the best defined and uniform known sequence of such binaries.

Submitted to: Astronomical Journal

For preprints, contact rucinski@astro.utoronto.ca
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0006236

The stars near the centre of supergiant shell LMC 4: Further constraints on triggering scenarios

Jochen M. Braun, Klaas S. de Boer, Martin Altmann

Sternwarte der Universität Bonn, Auf dem Hügel 71, D–53121 Bonn, Germany

The huge supergiant shell (SGS) LMC 4, next to the giant H ii region 30 Doradus the second most impressive feature of the Large Magellanic Cloud (LMC), has been the subject of many studies and was used to test various models of star forming mechanisms.

In this paper we present B, V photometry of the very centre of this SGS. The new data yield an age of 11 (2) Myr and a B −V excess of 0.10 (3) mag. Additionally, the two clusters near the geometric centre are too old to be related to the formation of LMC 4: HS 343 being ∼ 0.1 Gyr and KMHK 1000 being ∼ 0.3 Gyr of age. This and existing photometries support that the young stellar population covering the entire inner region is nearly coeval.

We compare these findings with predictions of models proposed for the creation mechanism of SGSs. We conclude that the large-scale trigger, necessary to explain the observations, comes from hydrodynamic interaction of the galactic halo with gas in the leading edge of the moving LMC.

Submitted to: Astronomy & Astrophysics

For preprints, contact jbraun@astro.uni-bonn.de
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0006060
or via ftp from ftp.astro.uni-bonn.de/pub/jbraun/p2 lmc4c/ (incl. photometry tables)
For additional information see http://www.astro.uni-bonn.de/˜ jbraun/phdt lmc4c.html

The Metallicity Distribution Function of Red Giants in the LMC^{1}

A.A. Cole^{2}, T.A. Smecker-Hane^{3}, & J.S. Gallagher III^{4}

1 Based on observations obtained at Cerro Tololo Inter-American Observatory. 2 University of Massachusetts, Amherst, USA 3 University of California, Irvine, USA 4 University of Wisconsin–Madison, USA

We report new metallicity determinations for 39 red giants in a 220 square arcminute region, 1.^{∘}8 southwest of the bar of the Large Magellanic Cloud. These abundance measurements are based on spectroscopy of the Ca ii infrared triplet. We have carefully considered the effects of abundance ratios, the physics of Ca ii line formation, the variation of red clump magnitude, and the contamination by foreground stars in our abundance analyses. The metallicity distribution function (MDF) shows a strong peak at [Fe/H] = −0.57 ± 0.04; a tail to abundances at least as low as [Fe/H] ≈ −1.6 brings the average abundance down to [Fe/H] = −0.64 ± 0.02. Half the red giants in our field fall within the range −0.83 ≤ [Fe/H] ≤ −0.41. The MDF appears to be truncated at [Fe/H] ≈ −0.25; the exact value of the maximum abundance is subject to ∼0.1 dex uncertainty in the calibration of the Ca ii IR triplet for young, metal-rich stars. We find a striking contrast in the shape of the MDF below [Fe/H] ≤ −1 between our inner disk field and the distant outer field studied by Olszewski (1993): red giants deficient by more than a factor of ten in heavy elements relative to the Sun are extremely scarce in the inner disk of the LMC. Our field star sample does not reproduce the full MDF of the LMC star clusters, but seems similar to that of the intermediate-age (1–3 Gyr) clusters. We have also obtained abundance estimates using Strömgren photometry for ≈ 10^{3} red giants in the same field. Photometry is the only practical way to measure abundances for the large numbers of stars necessary to lift age-metallicity degeneracy from our high-precision color-magnitude diagrams. The Strömgren measurements, which are sensitive to a combination of cyanogen and iron lines, correlate well with the Ca ii measurements, but a metallicity-dependent offset is found. The offset may be due either to variations in the elemental abundance ratios due to galactic chemical evolution, or to a metal-dependent mixing mechanism in RGB stars. An empirical relation between photometric and spectroscopic abundance estimates is derived. This will allow photometric abundance measurements to be placed on a consistent metallicity scale with spectroscopic metallicities, for very large numbers of stars.

Accepted by: The Astronomical Journal (October 2000)

For preprints, contact cole@condor.astro.umass.edu
Also available from the URL http://xxx.lanl.gov/ps/astro-ph/0006327

The Rotation Curve of the Large Magellanic Cloud and the Implications for Microlensing

David R. Alves^{1} and Cailin A. Nelson^{2}

1 Space Telescope Science Institute, Baltimore, MD 21218, USA 2 Department of Physics, University of California, Berkeley, CA 94720, USA

The rotation of the disk of the Large Magellanic Cloud (LMC) is derived from the radial velocities of 422 carbon stars (Kunkel, Irwin, & Demers 1997, A&AS, 122, 463). New aspects of this analysis include the propagation of uncertainties in the LMC proper motion with a Monte Carlo, and a selfconsistent modeling of the rotation curve and disk kinematics. The LMC rotation curve reaches a maximum circular velocity of 72 ± 7 km s^{−1} at R = 4.0 kpc and then declines. The rotation curve is well fit by a truncated, finite-thickness exponential disk model with no dark halo, implying a total mass of 5.3 ± 1.0 × 10^{9} M_{⊙}. The velocity dispersion in concentric radial bins from R = 0.5 to 5.6 kpc decreases from 22 to 15 km s^{−1}, then increases to ∼20 km s^{−1} at larger radii. Constant-thickness disk models in virial equilibrium cannot be reconciled with the data even if the effects of LMC or Galactic dark halos are included. If the disk is virialized, the scale height rises from h = 0.3 to 1.6 kpc over the range of R = 0.5 to 5.6 kpc. Thus the LMC disk is flared. We model the velocity dispersion at large radii (R > 6 kpc) as a maximal flared disk under the influence of the Galactic dark halo, which favors a mean density for the latter of \overline{ρ} ∼ 2.5 × 10^{−4} M_{⊙} pc^{−3} at the LMC distance.

LMC stellar kinematics play an important role in elucidating the nature of MACHOs, a dark population inferred from LMC microlensing. We have constructed a truncated and flared maximal disk model for the LMC which is kinematically based. Our model does not include a nonvirialized component such as tidal debris. The instantaneous probability of microlensing from LMC stars in our model is τ < 1.0 × 10^{−8} · sec^{2}i, where i is the disk inclination. Our upper limit on the self-lensing optical depth is in good agreement with that obtained from less sophisticated models, and is an order of magnitude too small to account for the MACHO microlensing signal.

Accepted by: The Astrophysical Journal For preprints, contact alves@stsci.edu

Also available from the URL http://arXiv.org/abs/astro-ph/?006018

Meeting Announcement

Modes of Star Formation and the Origin of Field Populations

Date: October 9 to 13, 2000

Venue: Max Planck Institute for Astronomy, Heidelberg, Germany

URL: http://www.mpia-hd.mpg.de/~modes/
Contact: Eva K. Grebel, modes@www.mpia-hd.mpg.de

Scientific Rationale: The purpose of this meeting is to identify and discuss the different modes of star formation in the local universe and how these contribute to the origin and evolution of field stellar populations in galaxies.

Topics to be discussed include: Which processes govern the differing star formation efficiencies and modes of star formation? To what extent does galaxy environment play a role? What are the spatial scales over which star formation typically occurs, and how do they depend on galaxy mass, type, rotation patterns, gas content, and location? Over what time scales does star formation occur within a given region? Are external trigger mechanisms required? Do field populations in galaxies generally originate in gradually dispersing associations and from dissolved open clusters? Can star formation occur in ”isolation” in the field, and if so what range of masses can be formed this way, and how common is this process? How does the initial mass function depend on environment and formation conditions?

The meeting consists of a mixture of invited and contributed talks as well as discussion sessions. Poster presenters may have up to 10 min for an oral summary of their contribution.

Speakers include: Tim Beers, Anthony Brown, Sofia Feltzing, Annette Ferguson, Yasuo Fukui, Reinhard Genzel, Thomas Henning, Paul Hodge, Deidre Hunter, Charles Lada, Mordecai Mac Low, Thomas Preibisch, Harri Vanhala

SOC: Wolfgang Brandner, Andi Burkert, Bruce Elmegreen, Jay Gallagher, Eva Grebel (Chair), Rob Kennicutt, Pavel Kroupa, Harold Yorke

Proceedings: The proceedings will be published in the Astronomical Society of the Pacific (ASP) Conference Series. The cost will be included in the registration fee.

Registration: The early registration deadline is July 15, 2000. Due to space and logistic limitations the number of attendees will be limited to ∼ 50 people. A registration form is available from the above URL. Participants will be selected on a first-come, first-served basis and based on how well their proposed contribution fits in with the topic of the meeting.

Editors’ Note:

The following meeting announcement is unrelated to Magellanic Clouds research. We agreed to include it anyway since it may be of general interest to some of the readers of MCNews who plan to attend the IAU General Assembly in August.

Networking Lunch for Women Astronomers During the IAU General Assembly in Manchester

The Royal Astronomical Society’s Committee for Women in Astronomy and Geophysics intend to organise an informal lunchtime meeting on Tuesday 8th August. The aim of this meeting is to provide an informal forum for discussion of the various issues which particularly affect women astronomers.

We hope to invite two or three women astronomers to say a few words on their own experiences to provide a focus for general discussion and we will write up a report of the meeting for Astronomy and Geophysics magazine.

We will provide a light buffet lunch although as our budget is limited, we encourage anyone who has already pre-booked a packed lunch to take it along. The venue is yet to be decided but will be on the main Manchester University site. The choice of venue will depend on the expected number of participants.

Any GA24 participants (including men and non-astronomer partners) who would be interested in attending this meeting are invited to contact the organiser:

Dr. Myfanwy Bryce Jodrell Bank Observatory Email: mbryce@ast.man.ac.uk Phone: +44 1477 572624 FAX: +44 1477 571618